3.2 moles of ammonia (NH3) is formed in the following reaction. How many grams of hydrogen reacted? 2 sig figs in your answer.
3H2 + N2 --> 2NH3
1
Expert's answer
2020-04-28T12:20:33-0400

"3H_2 + N_2 \to 2NH_3"

As per stochiometry

"2" moles of "NH_3" require "3" moles of "H_2."

"3.2" moles of "NH_3" will produce "\frac{3}{2}\times 3.2=4.8" moles of "H_2" .

Molar mass of "H_2" is "1\ g."

Mass of "H_2" required"=1\times 4.8=4.8\ g."